OpenStudy (anonymous):

help please? how to simplify the following radicals or equations. can someone tell me how, not just the answer, too? 4/sqrt 12x^3 (3sqrt11+5)^2 (-3sqrt14x^3)/-sqrt21x 2sqrtx = sqrt6x + 5 okay, this is the last one... 2x = sqrt5x+6 please and thank you!!!

OpenStudy (mathmale):

\[\frac{ 4 }{ \sqrt{12x^3} }\]

OpenStudy (mathmale):

cannot be simplified to the point where it has no radicals. However, you can "rationalize the denominator," which means to eliminate the radical from the denominator (only).

OpenStudy (mathmale):

First, let's actually do it! Then, we can discuss what we did and why. Multiply both numerator and denominator of the following by Sqrt(12x):\[\frac{ 4 }{ \sqrt{12x^3} }\frac{ \sqrt{12x} }{\sqrt{12x}}\]

OpenStudy (mathmale):

Would you now do the necessary math?

OpenStudy (anonymous):

so 4sqrt12x^3/12x^3

OpenStudy (mathmale):

I agree with the ' 12 ' in the denominator, but not with the x^3. Mind reviewing this and seeing if you can fix it?

OpenStudy (mathmale):

Look at the denominator again: It includes \[\sqrt{12x^3}\sqrt{12x}\]

OpenStudy (mathmale):

we can re-write this as \[\sqrt{12^2}\sqrt{x^3*x}\]

OpenStudy (mathmale):

How would you simplify that?

OpenStudy (mathmale):

Hints:\[\sqrt{2^2}=2,~and~\sqrt{x^4}=x^2\]

OpenStudy (mathmale):

@quickrabbit: OpenStudy informs me that you are currently offline. Next time, would you please let me know if you're signing off. I would have been willing to lead you through the solution of this problem, had you stuck with it.